OpenStudy (anonymous):

What is the vertex for the graph of the quadratic function f(x)=1/4(x+2)^2-9

OpenStudy (anonymous):

(-2,-9) Same reason as your previous question.

OpenStudy (anonymous):

Ok please explain. Vertex form is y=a(y-h)^2+k

OpenStudy (anonymous):

(h,k)

OpenStudy (anonymous):

ok. now how do i work that out?

OpenStudy (anonymous):

it should be : y=a(x-h)^2+k

OpenStudy (anonymous):

in the formula foolformath gave f(x)=1/4(x+2)^2-9 y=a(x-h)^2+k you can now determine the center w/c is h,k

OpenStudy (anonymous):

ok...i see

OpenStudy (anonymous):

thank you both